# Multi-Source Temporal Attention Network for Precipitation Nowcasting

### v1

@@ 1 Introduction
In agriculture, the main focus of [anonymized company], precipitation nowcasting can contribute to combating climate change by reducing fuel consumption for unsuccessful trips to fields during unfavorable weather, optimizing timing of grain harvesting and drying to reduce energy use and dry matter loss, enhancing spray efficiency by dosage of products for weather conditions, and preventing product waste by scheduling plant protection treatments at optimal times.
@@ 2 Approach: multi-source temporal attention
The Danish Meteorological Institute (DMI) provides radar composite data that captures rainfall intensities at 10-minute intervals with a resolution of 500 meters per pixel  [^1] https://opendatadocs.dmi.govcloud.dk/Data/Radar_Data. Still, its range is limited due to its ground-based nature (cf. Fig. A.1). To provide a more comprehensive view of the atmospheric state, we propose to complement the radar data with additional data from geostationary EUMETSAT satellites [^2] European Organisation for the Exploitation of Meteorological Satellites https://user.eumetsat.int/data/satellites/meteosat-second-generation covering broader regions, but with lower resolution. We obtain physical properties from GFS satellite imagery [^3] Global Forecast System managed by National Oceanic and Atmospheric Administration (NOAA), United States https://www.emc.ncep.noaa.gov/emc/pages/numerical_forecast_systems/gfs.php, which provides forecasts with a spatial resolution of 0.25 degrees and hourly temporal resolution for the first five days of the forecast in addition to the current state of the atmosphere with derived physical measurements. We process this data to generate sequences of patches of size, resolution, and context optimized for each source to fit GPU memory, and use a sliding window with blackout periods to prevent data leakage (details in Appendix B).
Architecture Recurrent networks suffer from poor computational efficiency, motivating us to leverage the Temporal Attention Unit (TAU) [tau] which features a spatial encoder and a decoder for intra-frame features, with temporal modules stacked in between to extract time-dependent features. A residual connection between the encoder and decoder preserves spatial information. The temporal module, built for parallel processing, uses depth-wise convolutions, dilated depth-wise convolutions, and 1×1 convolutions to address long-range dependencies. Pooling across the spatial dimension and fully-connected layers across the temporal dimension allow to learn temporal variations. The architecture in [tau] has fixed number of timesteps and channels. To handle data sources with different timesteps, sizes, and resolutions, our encoder standardizes all inputs independently to the same resolution and size before feeding them into the temporal module. Additionally, our decoder includes a residual connection for each resolution and produces a single timestep for the specified lead time with channels to represent various rain intensities (Fig. 1).
@@
Loss function We employ cross-entropy loss for probabilistic forecasts with comprehensive information [metnet, metnet2, metnet3]. We design classes of different rain intensities to have narrower ranges for lower precipitation values more frequently observed, while still encompassing rarer instances of heavy rain. Cross-entropy loss treats mispredictions equally, regardless of class. We instead propose weighing per-pixel loss by the difference between the indices of the target class and of the predicted class.
@@ 3 Experiments and results
A sample forecast, ground truth, GFS and Harmonie forecasts is shown in Figure 3. As lead time increases, uncertainty grows, resulting in more blurry precipitation patterns. Still, our model identifies the high precipitation forming in later lead times, unlike GFS. While our forecast appears more blurred than Harmonie, it actually improves accuracy by accounting for inherent uncertainty in predicting up to 8 hours ahead. Harmonie is noticeably shifted, which results in worse performance.
Figure 3: Sample visualization of ground truth, model prediction, and GFS and Harmonie forecasts. Even though our model provides predictions at 10-minute intervals, hourly intervals are shown.
@@ 4 Discussion and conclusion
We present a precipitation nowcasting model for Denmark that surpasses existing operational systems for up to 8 hours by leveraging multiple data sources, an advanced spatio-temporal architecture, optimized training with quality maps, and dynamic thresholds. Future work primarily involves expanding coverage to Europe radar from OPERA. Incorporating sparse observations from a network of >3,500 European weather stations may correct radar observations with surface rain gauge measurements.

### v2

@@ 1 Introduction
In agriculture, the main focus of Cordulus, precipitation nowcasting can contribute to combating climate change by reducing fuel consumption for unsuccessful trips to fields during unfavorable weather, optimizing timing of grain harvesting and drying to reduce energy use and dry matter loss, enhancing spray efficiency by dosage of products for weather conditions, and preventing product waste by scheduling plant protection treatments at optimal times.
@@ 2 Approach: multi-source temporal attention
The Danish Meteorological Institute (DMI) provides radar composite data that captures rainfall intensities at 10-minute intervals with a resolution of 500 meters per pixel  [^1] https://opendatadocs.dmi.govcloud.dk/Data/Radar_Data. Still, its range is limited due to its ground-based nature (cf. Fig. A.1). To provide a more comprehensive view of the atmospheric state, we propose to complement the radar data with additional data from geostationary EUMETSAT satellites [^2] European Organisation for the Exploitation of Meteorological Satellites https://user.eumetsat.int/data/satellites/meteosat-second-generation covering broader regions, but with lower resolution. We obtain physical properties from GFS satellite imagery [^3] Global Forecast System managed by National Oceanic and Atmospheric Administration (NOAA), United States https://www.emc.ncep.noaa.gov/emc/pages/numerical_forecast_systems/gfs.php, which provides forecasts with a spatial resolution of 0.25 degrees and hourly temporal resolution for the first five days of the forecast in addition to the current state of the atmosphere with derived physical measurements. We process this data, spanning from January 2022 to May 2024, to generate sequences of patches of size, resolution, and context optimized for each source to fit GPU memory, and use a sliding window with blackout periods to prevent data leakage (details in Appendix B).
Architecture Recurrent networks suffer from poor computational efficiency, motivating us to leverage the Temporal Attention Unit (TAU) [tau] which features a spatial encoder and a decoder for intra-frame features, with temporal modules stacked in between to extract time-dependent features. A residual connection between the encoder and decoder preserves spatial information. The temporal module, built for parallel processing, uses depth-wise convolutions, dilated depth-wise convolutions, and 1×1 convolutions to address long-range dependencies. Pooling across the spatial dimension and fully-connected layers across the temporal dimension allow to learn temporal variations. The architecture in [tau] has fixed number of timesteps and channels. To handle data sources with different timesteps, sizes, and resolutions, our encoder standardizes all inputs independently to the same resolution and size before feeding them into the temporal module. Additionally, our decoder includes a residual connection for each resolution and produces a single timestep for the specified lead time with channels to represent various rain intensities (Fig. 1). Instead of a continuous map, we predict probabilities in intensity bins to highlight both common light and rare heavy rainfall.
@@
Loss function We employ cross-entropy loss for probabilistic forecasts to capture comprehensive information [metnet, metnet2, metnet3]. We design classes of different rain intensities to have narrower ranges for lower precipitation values more frequently observed, while still encompassing rarer instances of heavy rain. Cross-entropy loss treats mispredictions equally, regardless of class. We instead propose weighing per-pixel loss by the difference between the indices of the target class and of the predicted class.
@@ 3 Experiments and results
A sample forecast, ground truth, GFS, Harmonie, and PySTEPS forecasts is shown in Figure 3. As lead time increases, uncertainty grows, resulting in more blurry precipitation patterns. Still, our model identifies the high precipitation forming in later lead times, unlike GFS. While our forecast appears more blurred than Harmonie, it actually improves accuracy by accounting for inherent uncertainty in predicting up to 8 hours ahead. Harmonie is noticeably shifted, which results in worse performance. PySTEPS accurately predicts the earlier lead times but quickly becomes ineffective due to the constant motion and intensity assumption.
Figure 3: Sample ground truth, model prediction, GFS, Harmonie, and PySTEPS forecasts. Even though our model provides predictions at 10-minute intervals, hourly intervals are shown.
@@ 4 Discussion and conclusion
We present a precipitation nowcasting model for Denmark that surpasses existing operational systems for up to 8 hours by leveraging multiple data sources, an advanced spatio-temporal architecture, optimized training with quality maps, and dynamic thresholds. Future work primarily involves expanding coverage to Europe radar from OPERA. Incorporating sparse observations from Cordulus’ >4,000 European weather stations may correct radar observations with surface rain gauge measurements.
Acknowledgments and Disclosure of Funding
This work is partly funded by the Innovation Fund Denmark (IFD) under File No. 2052-00064B.